If two operator do not commute what does it mean
1
Expert's answer
2019-06-25T09:10:29-0400

From the definition

"[A, B] = AB - BA \neq 0"

i.e. there exists such state

"A(B|\psi\rangle) \neq B(A|\psi\rangle)"

Moreover, if A is hermitian, there must exists an eigenstate of A, that is not an eigenstate of B. Proof by contradiction: let any eigenstate of A with the eigenvalue a is an eigenstate of B with the eigenvalue b. Then


"AB|\psi\rangle = Ab|\psi\rangle = ab|\psi\rangle = BA|\psi\rangle"

and due to the fullness of the set of eigenvectors of a hermitian operator this equality is true for any state - wrong.

Hence there exists a state, in which both quantities A and B cannot be determined simultaneously.


Heisenberg's uncertainty principle: consider displacement

"\Delta A = A - \langle A \rangle""\langle\Delta A^2 \rangle= \langle (A - \langle A \rangle)^2\rangle = \langle A^2 \rangle - \langle A \rangle^2"

Then in any state


"\langle \Delta A^2 \rangle \langle \Delta B^2 \rangle \geq\frac{\langle C \rangle^2}{4}"

where we define

"i C \equiv [A,B]"


Hence C is hermitian:


"-iC^\dagger = [A,B]^\dagger = (AB-BA)^\dagger = B^\dagger A^\dagger - A^\dagger B^\dagger = [B,A] = -[A,B] = -i C"

Proof: consider

"\langle (A -i\alpha B)\psi | (A - i\alpha B) \psi \rangle = \langle \psi | (A +i\alpha B) (A -i\alpha B) |\psi \rangle \geq 0""\langle \Delta A^2 - i \alpha [\Delta A, \Delta B] +\alpha^2\Delta B^2 \rangle \geq 0"

using

"[\Delta A, \Delta B] = [A,B] = iC"

and the condition that the discriminant of the inequality with respect to alpha is non-negative:

"\langle C \rangle ^2 - 4 \langle \Delta A^2 \rangle \langle \Delta B^2 \rangle \geq 0""\langle \Delta A^2 \rangle \langle \Delta B^2 \rangle \geq\frac{\langle C \rangle^2}{4}"

For example, it quantum mechanics

"[x, p] = i\hbar"

Here A = x, B = p and C = h, hence

"\langle \Delta x^2 \rangle \langle \Delta p^2 \rangle \geq\frac{\hbar^2}{4}"

in any state.